(1 point) Let f(x) = ? Compute (5+4x)/x f(x) = f(x) = f(1) = f(1) = f(1) = f(1) = f(iv) (1) = f (1) = f(iv)(x) = f(V)(x)
Answer #1

Salmo Fas= 5* 5+42 compute f(1) = 1 9 f\x) = F(x) = fles = 5] {(19 = [10 10 Io ) = 1-30 24 -) fle (1) = [120] 120 = f (*) -